on a standarized test with a mean of 500 and and a stadard deviation of 100 what would be the cutoff score for selecting the top 10% of the applicants
Answer by stanbon(57285)

on a standarized test with a mean of 500 and and a standard deviation of 100 what would be the cutoff score for selecting the top 10% of the applicants
---
Find the z-value with a left-tail of 90%
---
invNOrm(0.90) = 1.2816
----
cutoff = z*s + u
cutoff = 1.281556*100 + 500 = 628.1556
========================================
Cheers,
Stan H.
=============